# Understanding Ethereum Mempool Security under Asymmetric DoS by Symbolic Fuzzing

### v2

@@ Abstract
In blockchains, mempool controls transaction flow before consensus, denial of whose service hurts the health and security of blockchain networks. This paper presents mpfuzz, the first mempool fuzzer to find asymmetric DoS bugs by symbolically exploring mempool state space and optimistically estimating the promisingness an intermediate state is in reaching bug oracles. Compared to the baseline blockchain fuzzers, mpfuzz achieves a >100×>100$\times$> 100 × speedup in finding known DETER exploits. Running mpfuzz on six major Ethereum clients leads to the discovering of new mempool vulnerabilities, which exhibit a wide variety of sophisticated patterns including stealthy mempool eviction and mempool locking. Rule-based mitigation schemes are proposed against newly discovered vulnerabilities.
@@ 1 Introduction
In Ethereum, a mempool buffers unconfirmed transactions from web3 users before they are included in the next blocks. Mempool provides the essential functionality to bridge the gap between varying rates of submitted transactions and rates of produced blocks, despite what types of transactions it serves, whether public or private transactions. As shown in recent studies [22], denying a mempool service can force the blockchain to produce blocks of low or even zero (Gas) utilization, undermining validators’ incentives and shrinking the blockchain networks in the long run, re-introducing the $51\%$ attacks. Besides, a denied mempool service can drop normal transactions, cutting millions of web3 users off the blockchain and failing the decentralized applications relying on real-time blockchain access.
@@
Problem: Spamming the mempool to deny its service has been studied for long [17, 9, 27, 12]. Early designs by sending spam transactions at high fees burden attackers with high costs and are of limited practicality. What poses a real threat is Asymmetric DeniAl of Mempool Service, coined by ADAMS, in which the mempool service is denied at an asymmetrically low cost. That is, the attack costs in terms of the fees of adversarial transactions are significantly lower than those of normal transactions victimized by the denied mempool. In the existing literature, DETER [22] is the first ADAMS attack which works by sending invalid transactions to directly evict normal transactions in the mempool. MemPurge [29] is a similar mempool attack that finds a way to send overdraft transactions into Geth’s pending-transaction pool and cause eviction there. These known attacks are easy to detect (i.e., following the same direct-eviction pattern), and in fact, the DETER bugs reported in 2021 have been successfully fixed in all major Ethereum clients as of Fall 2023, including Geth, Nethermind, Erigon, and Besu. Given this state of affairs, we pose the following research question: Are there new ADAMS vulnerabilities in the latest Ethereum clients already patched against direct-eviction based attacks?
@@
This work takes an automatic and systematic approach to discovering ADAMS vulnerabilities, unlike the existing DETER bugs that are manually found. Fuzzing mempool implementations is promising but also poses unique challenges: Unlike the consensus implementation that reads only valid confirmed transactions, the mempool residing pre-consensus needs to handle various unconfirmed transactions, imposing a much larger input space for the fuzzer. For instance, a mempool can receive invalid transactions under legitimate causes, [^1] For instance, future transactions can be caused by out-of-order information propagation in Ethereum. and factors such as fees are key in determining transaction admission outcomes. Existing blockchain fuzzers including Fluffy [30], Loki [23] and Tyr [19] all focus on fuzzing consensus implementation and don’t explore the extra transaction space required by mempool fuzzing. As a result, the naive application of a consensus fuzzer to fuzz mempool misses detecting the DETER bugs as evaluated in § 4.1, let alone discover more sophisticated new ADAMS bugs.
@@
Proposed methods: To enable efficient mempool fuzzing, our key observation is that real-world mempool implementations admit transactions based on abstract “symbols”, such as pending and future transactions, instead of concrete value. Sending multiple concrete transactions generated under the same symbol triggers the same mempool behavior and should be avoided. Mutating one transaction under one symbol during fuzzing suffices to cover mempool behavior.
We propose symbolic stateful mempool fuzzing or mpfuzz. We first extract seven symbols from Ethereum semantics to cover future, parent, child, overdraft, latent overdraft, and replacement transactions. Then, mpfuzz generates a new transaction by symbol, sends the input to execute on a target mempool implementation, observes the end state of the mempool, and uses the symbolic state coverage as feedback to guide the next round of mutation. By this means, mpfuzz searches the symbolic mempool state space for the one matching the bug oracles induced from ADAMS. To further speed up the search in a limited time, we prioritize “promising” states by optimistically estimating the costs of unconfirmed transactions whose validity is subject to change in the future.
@@
Found attacks: We run mpfuzz on six Ethereum execution-layer clients including Geth [2], Nethermind [5], Erigon [1], Besu [3], Reth [7], and OpenEthereum [6]. Compared to the baseline fuzzers, mpfuzz can find exploits faster by more than two orders of magnitude. On Ethereum clients of historical versions, mpfuzz can rediscover known DETER bugs. On the latest Ethereum clients where DETER bugs are fixed, mpfuzz can find new ADAMS attacks, described next.
@@
We summarize the newly discovered ADAMS exploits into several patterns: 1) Indirect eviction by valid-turned-invalid transactions. Unlike the direct-eviction pattern used in DETER, the indirect-eviction attack works more stealthily in two steps: The attacker first sends normal-looking transactions to evict victim transactions from the mempool and then sends another set of transactions to turn the admitted normal-looking transactions into invalid ones, bringing down the cost. 2) Locking mempool, which is to occupy the mempool to decline subsequent victim transactions. Mempool locking does not need to evict existing transactions from the mempool as the eviction-based DETER attacks do. 3) Adaptive attack strategies where the attack composes adversarial transactions in an adaptive way to the specific policies and implementations of the mempool under test. Example strategies include composing transactions of multiple patterns in one attack, re-sending evicted transactions to evict “reversible” mempools, locking mempool patched against turning, etc.
We propose rule-based mitigation schemes against some of our discovered bugs and identify design challenges in mitigating others, which is left for future works.
@@
All newly found ADAMS bugs are reported to the Ethereum Foundation with $7$ bugs confirmed and $3$ fixed in recent client releases [4]. Bug reporting is documented on the webpage [13] which also includes demos of mpfuzz on six Ethereum clients. mpfuzz will be open-sourced to facilitate vulnerability discovery in more and future clients.
@@
$\bullet$ New fuzzing problem: This paper is the first to formulate the mempool-fuzzing problem to automatically find asymmetric DoS vulnerabilities as bug oracles. Fuzzing mempools poses new unique challenges that existing blockchain fuzzers don’t address and entails a larger search space including invalid transactions and varying fees.
$\bullet$ Symbolic fuzzing tool: The paper presents the design and implementation of mpfuzz, a symbolic stateful mempool fuzzer. Given a mempool implementation, mpfuzz searches the mempool state space using the feedback of symbolic state coverage and promising-ness to reach bug oracles. With a Python prototype and runs on real Ethereum clients, mpfuzz achieves a >100×>100$\times$> 100 × speedup in finding known DETER exploits compared to baselines.
$\bullet$ New mempool vulnerabilities: mpfuzz discovers new asymmetric-DoS vulnerabilities in six major Ethereum clients in the mainnet. By evaluation under real transaction workloads, all found attacks achieve $84.6-99.6\%$ success rates and low costs such as adversarial transaction fees 100×100$\times$100 × lower than the victim transaction fees.
@@ 2 Related Works
Loki [23] is a stateful fuzzer to find consensus bugs causing crashes in blockchain network stacks. Specifically, Loki runs as a fuzzer node interacting with a tested blockchain node through sending and receiving network messages. 1) The test oracle in Loki is whether the tested blockchain node crashes. 2) The fuzzer node uses a learned model of blockchain network protocol to generate the next message of complying format, in which the content is mutated by randomly choosing integers and bit-flipping strings. In other words, the mutation in Loki is unaware of application level semantics. For instance, if the message is about propagating an Ethereum transaction, the transaction’s nonce, Ether amount, and Gas price are all randomly chosen; the sender, receiver, and data are bit flipped. 3) Loki records messages sent and received from the test node. It uses them as the state. New states receive positive feedback.
@@
For mempool DoS, a basic attack is by sending spam transactions with high fees to evict victim transactions of normal fees [17, 9, 27, 12], which incur high costs to attackers. DETER [22] is the first asymmetric DoS on Ethereum mempool where the adversary sends invalid transactions of high fees (e.g., future transactions or latent overdrafts) to evict victim transactions of normal fees. Concurrent to this work is MemPurge [29] posted online in June 2023; in it, the attacker reconnects her adversarial future transactions and makes them latent-overdraft transactions. As we tested on the latest Ethereum clients in July 2023, such as Geth $v1.11.4$, the DETER bugs have been fixed. Both DETER and MemPurge are manually discovered.
3 Threat Model
In the threat model, an attacker aims at disabling the mempools on one or multiple nodes and causing the disruption of the transaction propagation service. This can lead to damage, such as delayed or dropped transaction inclusion and low-utilization blocks (even empty ones). The damage can benefit certain attackers in practice, such as competing RPC services, MEV bots, or block validators. Besides, the attacker aims at causing the above damage at asymmetrically low costs, that is, the fees of adversarial transactions included in the blockchain are lower than the fees of victim transactions outside the mempool, as will be presented.
The attacker has the same capability as in the DETER work [22], where the attacker joins an Ethereum network, discovers critical nodes (e.g., backend nodes of a RPC service, a node hosting MEV bot, or target block validators), and sends adversarial transactions.
@@ 3.1 Notations
Annotation
Meaning
Symbol
Meaning
$m$
Mempool length
$ops$
Tx sequence
$st$
State of residential txs in mempool
$dc$
History of declined txs from mempool
$A$
Adversarial tx sender
$B$
Benign tx sender
$tx\begin{bmatrix}{s}&{v}\\ {n}&{f}\end{bmatrix}$
A transaction from sender $s$, of nonce $n$, transferring Ether value $v$, and with Gas Price $f$
@@
States: We recognize two mempool-related states in Ethereum, the collection of transactions stored in mempool $st$, and the collection of declined transactions $dc$.
@@
An arriving transaction can be admitted with eviction, admitted without eviction, or declined by a mempool. The three operations are defined as follows. 1) An arriving transaction $tx_{i}$ is admitted with evicting a transaction $tx_{i}\textquoteright$ ’ from the mempool, if $st_{i+1}=st_{i}\setminus{}tx_{i}^{\prime}\cup{}tx_{i},dc_{i+1}=dc_{i}$. 2) $tx_{i}$ is admitted without evicting any transaction in the mempool, if $st_{i+1}=st_{i}\cup{}tx_{i},dc_{i+1}=dc_{i}$. 3) $tx_{i}$ is declined and does not enter the mempool, if $st_{i+1}=st_{i},dc_{i+1}=dc_{i}\cup{}tx_{i}$.
3.2 Bug Oracles
Suppose a mempool of initial state $\langle{}st_{0},dc_{0}\rangle{}$ receives a sequence of transactions, denoted by $ops$. Here, we simplify the problem by not considering arriving blocks during fuzzing. We define two bug oracles: eviction-based attacks and declining-based locking attacks.

### v3

# Understanding Ethereum Mempool Security under Asymmetric DoS by Symbolized Stateful Fuzzing

@@ Abstract
In blockchains, mempool controls transaction flow before consensus, denial of whose service hurts the health and security of blockchain networks. This paper presents mpfuzz, the first mempool fuzzer to find asymmetric DoS bugs by exploring the space of symbolized mempool states and optimistically estimating the promisingness of an intermediate state in reaching bug oracles. Compared to the baseline blockchain fuzzers, mpfuzz achieves a >100×>100$\times$> 100 × speedup in finding known DETER exploits. Running mpfuzz on major Ethereum clients leads to discovering new mempool vulnerabilities, which exhibit a wide variety of sophisticated patterns, including stealthy mempool eviction and mempool locking. Rule-based mitigation schemes are proposed against all newly discovered vulnerabilities.
@@ 1 Introduction
In Ethereum, a mempool buffers unconfirmed transactions from web3 users before they are included in the next blocks. Mempool provides the essential functionality to bridge the gap between varying rates of submitted transactions and rates of produced blocks, regardless of public or private transactions it serves. As shown in recent studies [28], denying a mempool service can force the blockchain to produce blocks of low or even zero (Gas) utilization, undermining validators’ incentives and shrinking the blockchain networks in the long run, re-introducing the $51\%$ attacks. Besides, a denied mempool service can prevent normal transactions from block inclusion, cutting millions of web3 users off the blockchain and failing the DApps relying on real-time blockchain access.
@@
Problem: Spamming the mempool to deny its service has been studied for long [23, 15, 33, 18]. Early designs by sending spam transactions at high prices burden attackers with high costs and are of limited practicality. What poses a real threat is Asymmetric DeniAl of Mempool Service, coined by ADAMS, in which the mempool service is denied at an asymmetrically low cost. That is, the attack costs, in terms of the fees of adversarial transactions, are significantly lower than those of normal transactions victimized by the denied mempool. In the existing literature, DETER [28] is the first ADAMS attack, and it works by sending invalid transactions to directly evict normal transactions in the mempool. MemPurge [35] is a similar mempool attack that finds a way to send overdraft transactions into Geth’s pending transaction pool and causes eviction there. These known attacks are easy to detect (i.e., following the same direct-eviction pattern). In fact, the DETER bugs reported in 2021 have been successfully fixed in all major Ethereum clients as of Fall 2023, including Geth, Nethermind, Erigon, and Besu. Given this state of affairs, we pose the following research question: Are there new ADAMS vulnerabilities in the latest Ethereum clients already patched against direct-eviction based attacks?
@@
This work takes a systematic and semi-automated approach to discovering ADAMS vulnerabilities, unlike the existing DETER bugs that are manually found. Fuzzing mempool implementations is a promising approach but also poses unique challenges: Unlike the consensus implementation that reads only valid confirmed transactions, the mempool, which resides in the pre-consensus phase, needs to handle various unconfirmed transactions, imposing a much larger input space for the fuzzer. For instance, a mempool can receive invalid transactions under legitimate causes, [^1] For instance, future transactions can be caused by out-of-order information propagation in Ethereum. and factors such as fees or prices are key in determining transaction admission outcomes. Existing blockchain fuzzers including Fluffy [36], Loki [29] and Tyr [25] all focus on fuzzing consensus implementation and don’t explore the extra transaction space required by mempool fuzzing. As a result, directly re-purposing a consensus fuzzer to fuzz mempool would be unable to detect the DETER bugs as evaluated in Appendix § A, let alone discover more sophisticated new ADAMS bugs.
@@
Proposed methods: To efficiently fuzz mempools, our key observation is that real-world mempool implementation admits transactions based on abstract “symbols”, such as pending and future transactions, instead of concrete value. Thus, sending multiple transactions under one symbol would trigger the same mempool behavior repeatedly, and it suffices to explore just one transaction per symbol during fuzzing without losing the diversity of mempool behavior.
We propose symbolized-stateful mempool fuzzing or mpfuzz. To begin with, mpfuzz is set up and run in a three-step workflow: It is first manually set up against a size-reduced mempool under test, which induces a much smaller search space for fuzzing. Second, mpfuzz is iteratively run against the reduced mempool to discover short exploits. Third, the short exploits are manually extended to actual ones that are functional on real mempools and Ethereum clients. Internally, the design of mpfuzz is based on seven transaction symbols we design out of Ethereum semantics: future, parent, overdraft, latent overdraft, and replacement transactions. Under these symbols, in each iteration of fuzzing, mpfuzz explores one concrete transaction per symbol, sends the generated transaction sequence to the target mempool, observes the mempool end state, and extracts the feedback of symbolized state coverage and state promisingness in reaching bug oracles to guide the next round of fuzzing. In particular, mpfuzz employs a novel technique to evaluate state promisingness, that is, by optimistically estimating the costs of unconfirmed transactions whose validity is subject to change in the future.
@@
Found attacks: We run mpfuzz on six leading execution-layer Ethereum clients deployed on the mainnet’s public-transaction path (Geth [7], Nethermind [11], Erigon [5], Besu [10], Reth [13], and OpenEthereum [12]), three PBS builders (proposer-builder separation) on the mainnet’s private-transaction and bundle path (Flashbot $v1.11.5$ [6], EigenPhi [4] and bloXroute [1]), and three Ethereum-like clients (BSC $v1.3.8$ [2] deployed on Binance Smart Chain, go-opera $v1.1.3$ [9] on Fantom, and core-geth $v1.12.19$ [3] on Ethereum Classic). Compared to the baseline fuzzers, mpfuzz can find exploits faster by more than two orders of magnitude. On Ethereum clients of historical versions, mpfuzz can rediscover known DETER bugs. On the latest Ethereum clients where DETER bugs are fixed, mpfuzz can find new ADAMS attacks, described next.
@@
We summarize the newly discovered ADAMS exploits into several patterns: 1) Indirect eviction by valid-turned-invalid transactions. Unlike the direct-eviction pattern used in DETER, the indirect-eviction attack works more stealthily in two steps: The attacker first sends normal-looking transactions to evict victim transactions from the mempool and then sends another set of transactions to turn the admitted normal-looking transactions into invalid ones, bringing down the cost. 2) Locking mempool, which is to occupy the mempool to decline subsequent victim transactions. Mempool locking does not need to evict existing transactions from the mempool as the eviction-based DETER attacks do. 3) Adaptive attack strategies where the attack composes adversarial transactions in an adaptive way to the specific policies and implementations of the mempool under test. Example strategies include composing transactions of multiple patterns in one attack, re-sending evicted transactions to evict “reversible” mempools, locking mempool patched against turning, etc. We propose rule-based mitigation schemes against all newly discovered vulnerabilities.
@@
All newly found ADAMS bugs are reported to the developer communities, including Ethereum Foundation, BSC, Fantom, PBS builders, etc. with $15$ bugs confirmed and $4$ fixed in recent client releases [8]. Bug reporting is documented on the webpage [20], which also includes demos of mpfuzz on the tested Ethereum clients. mpfuzz will be open-sourced to facilitate vulnerability discovery in more and future clients.
@@
$\bullet$ New fuzzing problem: This paper is the first to formulate the mempool-fuzzing problem to automatically find asymmetric DoS vulnerabilities as bug oracles. Fuzzing mempools poses new unique challenges that existing blockchain fuzzers don’t address and entails a larger search space including invalid transactions and varying prices.
$\bullet$ New fuzzing method: The paper presents the design and implementation of mpfuzz, a symbolized-stateful mempool fuzzer. Given a mempool implementation, mpfuzz defines the search space by the mempool states covered under symbolized transaction sequences and efficiently searches this space using the feedback of symbolized state coverage and the promising-ness of an intermediate state in triggering bug oracles. With a Python prototype and runs on real Ethereum clients, mpfuzz achieves a >100×>100$\times$> 100 × speedup in finding known DETER exploits compared to baselines.
$\bullet$ New discovery of mempool vulnerabilities: mpfuzz discovers new asymmetric-DoS vulnerabilities in six major Ethereum clients in the mainnet. By evaluation under real transaction workloads, all found attacks achieve $84.6-99.6\%$ success rates and low costs such as adversarial transaction fees 100×100$\times$100 × lower than the victim transaction fees.
@@ 2 Related Works
Loki [29] is a stateful fuzzer to find consensus bugs causing crashes in blockchain network stacks. Specifically, Loki runs as a fuzzer node interacting with a tested blockchain node through sending and receiving network messages. 1) The test oracle in Loki is whether the tested blockchain node crashes. 2) The fuzzer node uses a learned model of blockchain network protocol to generate the next message of complying format, in which the content is mutated by randomly choosing integers and bit-flipping strings. That is, the mutation in Loki is unaware of application level semantics. For instance, if the message is about propagating an Ethereum transaction, the transaction’s nonce, Ether amount, and GasPrice are all randomly chosen; the sender, receiver, and data are bit flipped. 3) Loki records messages sent and received from the test node. It uses them as the state. New states receive positive feedback.
@@
For mempool DoS, a basic attack is by sending spam transactions with high fees to evict victim transactions of normal fees [23, 15, 33, 18], which incur high costs to attackers. DETER [28] is the first asymmetric DoS on Ethereum mempool where the adversary sends invalid transactions of high fees (e.g., future transactions or latent overdrafts) to evict victim transactions of normal fees. Concurrent to this work is MemPurge [35] posted online in June 2023; in it, the attacker reconnects her adversarial future transactions and makes them latent-overdraft transactions. The DETER bugs have been fixed on the latest Ethereum clients such as Geth $v1.11.4$, as we tested them in July 2023. Both DETER and MemPurge are manually discovered.
3 Background
Ethereum mempools: In Ethereum, users send their transactions to the blockchain network, which are propagated to reach validator nodes. On every blockchain node, unconfirmed transactions are buffered in the mempool before they are included in the next block or evicted by another transaction. In Ethereum 2.0, transactions are propagated in two fashions: public transactions are broadcast among all nodes, and private transactions are forwarded to selected nodes, more specifically, selected builders and proposers (as in PBS or proposer-builder separation). The mempool is present on both paths, and it is inside Ethereum clients such as Geth [7] and Nethermind [11] (handling public transactions), and Flashbot [6] (private transactions). On these clients, a mempool also serves many other downstream modules, including MEV searchers or bots (in PBS), Gas stations, RPC queries, etc.
Mempools on different nodes are operated independently and don’t need to synchronize as the consensus layer does. For instance, future transactions are not propagated, and the set of future transactions in the mempool on one node is different from another node.
The core mempool design is transaction admission: Given an initial state, whether and how the mempool admits an arriving transaction. In practice, example policies include those favoring admitting transactions of higher prices (and thus evicting or declining the ones of lower prices), transactions arriving earlier, certain transaction types (e.g., parent over child transactions), and valid transactions (over future transactions). See different types of Ethereum transactions in § 4.1.
Transactions and fees: In Ethereum, Gas measures the amount of computations caused by smart contract execution. When preparing for transaction $tx$, the sender needs to specify Gas and GasPrice; the former is the maximal amount of computations allowed for running smart contracts under $tx$, and the latter is the amount of Ether per Gas the sender is willing to pay. After $tx$ is included in a block, the actual amount of computations consumed by contract execution is denoted by GasUsed. The fees of transaction $tx$ are the product of GasUsed and GasPrice, that is, $tx.fee=\textit{GasUsed}*\textit{GasPrice}$.
After EIP-1559 [19], GasPrice is divided into two components: $BasePrice$ and $PriorityPrice$, that is, $\textit{GasPrice}=BasePrice+PriorityPrice$. Blockchain nodes follow the Ethereum protocol to derive $BasePrice$ from Gas utilization in recent blocks. The fees associated with the $BasePrice$ are burnt upon transaction inclusion. $PriorityPrice$ is set by the transaction sender.
4 Threat Model and Bug Oracle
In the threat model, an attacker controls one or few nodes to join an Ethereum network, discovers and neighbors critical nodes (e.g., top validators, backends of an RPC Service, MEV searcher) if necessary, and sends crafted adversarial transactions to their neighbors. The attacker in this work has the same networking capacity as in DETER [28].
The attacker’s goal is to disable the mempools on the critical or all nodes in an Etheurem network and cause damage such as dropped transactions in block inclusion, produced blocks of low or zero (Gas) utilization, and disabled other downstream services. In practice, this damage is beneficial to competing block validators, MEV searchers, and RPC services. Besides, the attacker aims to cause this damage at asymmetrically low costs, as will be described.
@@ 4.1 Notations
Symbol Meaning Symbol Meaning
$m$ Mempool length $ops$ Tx sequence
$st$ State of residential txs in mempool $dc$ History of declined txs from mempool
$A$ Adversarial tx sender $B$ Benign tx sender
$tx\begin{bmatrix}{s}&{v}\\ {n}&{f}\end{bmatrix}$ A transaction from sender $s$, of nonce $n$, transferring Ether value $v$, and with GasPrice $f$
@@
States and transitions: We recognize two mempool-related states in Ethereum, the collection of transactions stored in mempool $st$, and the collection of declined transactions $dc$.
@@
An arriving transaction can be admitted with eviction, admitted without eviction, or declined by a mempool. The three operations are defined as follows. 1) An arriving transaction $tx_{i}$ is admitted with evicting a transaction $tx_{i}^{\prime}$ from the mempool, if $st_{i+1}=st_{i}\setminus{}tx_{i}^{\prime}\cup{}tx_{i},dc_{i+1}=dc_{i}$. 2) $tx_{i}$ is admitted without evicting any transaction in the mempool, if $st_{i+1}=st_{i}\cup{}tx_{i},dc_{i+1}=dc_{i}$. 3) $tx_{i}$ is declined and does not enter the mempool, if $st_{i+1}=st_{i},dc_{i+1}=dc_{i}\cup{}tx_{i}$.
Figure 1: Overview of exploit discovery and evaluation workflow: 0) mpfuzz setup, 1) mempool reduction, 2) fuzzing on reduced mempool under test (MUT) to discover short exploits, and 3) exploit extension. The extended exploits are 4) evaluated on the actual mempool of the original size. Green means automated tasks, and gray requires manual effort.
4.2 Exploit Discovery Workflow
To set up the stage, we present the workflow overview in this work. The workflow is depicted in Figure 1. To discover vulnerabilities and exploits in an actual mempool, one has to 0) set up mpfuzz by various parameters (as will be introduced), 1) reduce the mempool from its original configuration to a much smaller version, named mempool under test (MUT). 2) mpfuzz is run on the reduced MUT deployed locally (with settings described in § 7.1). Fuzzing leads to the discovery of short exploits. 3) Short exploits are extended to the longer ones, that is, actual exploits. 4) The success of the actual exploit is evaluated on the actual mempools deployed in close-to-operational settings, such as using popular Ethereum testnets like Goerli (see § 6.3.1) or a network we set up locally running unmodified Ethereum clients (see § 4 and § D). Fuzzing a reduced mempool (MUT) instead of an original one is necessary to ensure the efficient execution of the mpfuzz, which entails many iterations of re-initiating and populating the mempool. We discuss and evaluate the possible false positives that could be introduced by fuzzing smaller MUT in § 7.2.
In this workflow, fully automated is Step 2), that is, the discovery of short exploits on a given MUT (colored green in Figure 1). Other steps colored in gray require manual efforts and are described below: In Step 0), mpfuzz setup entails setting parameters in bug oracles (e.g., $\epsilon$ and $\lambda$) and fixed symbols in guiding fuzzing. The setup is usually one-time. In Step 1), mempool reduction entails reconfiguring the mempool to a much smaller capacity (e.g., MUT length $m=6$ or $m=16$, compared to the original size like $m^{\prime}=6144$ as in Geth). In addition, policy-specific parameters in the actual mempool are tuned down proportionally. For instance, the Geth mempool buffers up to $py_{1}^{\prime}=1024$ future transactions, and when buffering more than $py^{\prime}_{3}=5120$ pending transactions, the mempool triggers a protective policy to limit the number of transactions sent from the same sender by $py_{2}^{\prime}=16$ In a reduced MUT of $m=6$, these parameters are reduced while retaining the same proportion, such as $py_{1}=1,py_{3}=5,py_{2}=2$. The mempool reduction requires expert understanding and manual efforts, and it is by nature the best effort. In Step 3), it first de-duplicates the raw short exploits found by mpfuzz. Then, for each distinct short exploit, it extends it by repeating the transaction-admission events to fit into the actual mempool, during which one may practice strategies like increasing nonce, switching sender accounts, and measuring normal transactions in the actual mempool. In exploit extension, one could also consider the factors not modeled in MUT or mpfuzz such as $BasePrice$, as seen in the case in § 6.4.
4.3 Bug Oracle: Definitions & Rationale
Definition 4.1 (Tx admission timeline)
In a transaction admission timeline, a mempool under test is initialized at state $\langle{}st_{0},dc_{0}=\varnothing\rangle{}$, receives a sequence of arriving transactions $ops$, and ends up with state $\langle{}st_{n},dc_{n}\rangle{}$. Then, a validator continually builds blocks by selecting and clearing transactions in the mempool until the mempool is empty. This transaction timeline is denoted by $\langle{}st_{0},dc_{0}=\varnothing\rangle{},ops\Rightarrow{}\langle{}st_{n},dc_{n}\rangle{}$.
The transaction admission timeline is simplified from the timeline that an operational mempool experiences where transaction arrival and block building can be interleaved. The simplification is intended to trade off the accuracy of fuzzing results for fuzzing efficiency as discussed in § 7.2.
We use two mempool-attack templates to define two bug oracles: an eviction-based DoS where adversarial transactions evict existing victim transactions in the mempool, and a locking-based DoS where existing adversarial transactions in the mempool decline arriving victim transactions.